assume there is uncertainty about the costs of disposing is the waste: there is 50 50 chance that they will bw $10000 of $30000. Discuss how this uncertainty effects the cost benefit calculation, if the government is risk neutral or very risk averse.
1
Expert's answer
2020-06-10T18:58:23-0400

The first step is to compute the expected cost. The expected cost will be:

"\text{Expected cost}=\dfrac{50}{100}\times 10,000+\dfrac{50}{100}\times 30,000"


"\text{Expected cost}=20,000"

This means that in cost benefit analysis, the government that is risk neutral would consider the cost of disposing to be $20,000. A risk averse government would take the cost of disposing waste to be more than $20,000.